Can you please help me solve this problem:
Three years ago, Rica was thrice as old as Dandy. Three years hence, she will be twice as old. Find their present ages.
: Can you please help me solve this problem:
Three years ago, Rica was thrice as old as Dandy. Three years hence, she will be twice as old. Find their present ages.

Answer by stanbon(18792)

Three years ago, Rica was thrice as old as Dandy. Three years hence, she will be twice as old. Find their present ages.
-----------------------
Let Rica's present age be "r". Let Dandy's present ate by "d".
-----------------------
EQUATIONS:
Three years ago: r-3 = 3(d-3)
Three years hence: r+3 = 2(d+3)
-----------------
Rearrange:
r - 3d = -6
r - 2d = 3
--------------
Subtract 1st from 2nd to get:
d = 9 (Dandy's age NOW)
-----------
Substitute into r-2d = 3
r -(2*9) = 3
r = 21 (Rita's age NOW)
================
Cheers,
Stan H.